Solve each equation.
184=98+g
v-4+5+17=56
28=c-44

Answer by Alan3354(69443)

184=98+g
v-4+5+17=56
28=c-44
------------
184=98+g
The idea is to isolate the variable, get it on one side of the equal by itself.
Subtract 98 from both sides.
86 = g
or, g = 86
That's all.
-------------
v-4+5+17=56
Collect terms
v + 18 = 56
Subtract 18
v = 38
-------
28=c-44
Add 44
72 = c
Time for a beer.